OpenStudy (anonymous):

nonlinear diff eq: dy/dx = y(y-1)/cos(x)

OpenStudy (fibonaccichick666):

can you figure out how to use separation of variables?

OpenStudy (anonymous):

1/cos(x) dx = 1/y(y-1) dy

OpenStudy (anonymous):

then integrate and solve for y

OpenStudy (anonymous):

yea i got it down to y^2/6(2y-3)=ln(sec(x)+tan(x) +c but i'm not sure how to solve for y from here

OpenStudy (fibonaccichick666):

oh, you don't solve for y anymore

OpenStudy (anonymous):

sorry: ln(y) + ln(y-1) = ln(sec(x)+tan(x) + c*

OpenStudy (fibonaccichick666):

solve for c

OpenStudy (fibonaccichick666):

but, your integral, are you sure it's correct?

OpenStudy (fibonaccichick666):

I'm getting very very close but not exactly

OpenStudy (fibonaccichick666):

@housey

OpenStudy (anonymous):

sorry just got it down to: ln(y-1) - ln(y) = ln|sec(x)+tan(x)| +c, Initial Condition: y(0) = 1/2

OpenStudy (fibonaccichick666):

uhm well I'm still not getting that exactly, let's check partial fractions on those ys

OpenStudy (anonymous):

so Integral(1/y^2-y) dy is the same as integral (-1/y) + integral (1/y-1) using partial fractions...?

OpenStudy (fibonaccichick666):

I got the opposite

OpenStudy (fibonaccichick666):

i got -1/y and 1/y-1

OpenStudy (fibonaccichick666):

sorry nvm dyslexic

OpenStudy (anonymous):

haha all good

OpenStudy (fibonaccichick666):

ok so now, we don't solve for y anymore, we just solve for C

OpenStudy (anonymous):

when i used the IC it came out as c = 0, see what u get

OpenStudy (fibonaccichick666):

or whatever you like to call the constant

OpenStudy (fibonaccichick666):

y(0)=1/2 right?

OpenStudy (anonymous):

yeah

OpenStudy (fibonaccichick666):

and cause it's 5 am y(0) is the same as y=0 x=0 right?

OpenStudy (anonymous):

so ln|-1/2| - ln|1/2| = ln|sec(0)+tan(0)|. meaning 0=ln(1) + c so 0= 0 + 0

OpenStudy (anonymous):

nah as in, when x=0, y= 1/2 :)

OpenStudy (anonymous):

love a bit of calculus at 5am

OpenStudy (fibonaccichick666):

ok sorry sped moment

OpenStudy (fibonaccichick666):

oh goodness

OpenStudy (anonymous):

no stress

OpenStudy (fibonaccichick666):

man I need some sleep haha

OpenStudy (fibonaccichick666):

watch your subs on the y

OpenStudy (fibonaccichick666):

should be ln|-2|+ln|1/(1/2-1)|

OpenStudy (fibonaccichick666):

ends up being ln|-2|+ln|2| right?

OpenStudy (anonymous):

not from where i'm sitting haha..: ln|0.5-1|-ln(05) ? so ln|-0.5| which equals ln(0.5) ?

OpenStudy (fibonaccichick666):

well ln|-2| but doesn't matter

OpenStudy (fibonaccichick666):

no it's not a minus

OpenStudy (fibonaccichick666):

where did you get the minus from

OpenStudy (fibonaccichick666):

we have ln|-1/y|+ln|1/(y-1)| no?

OpenStudy (anonymous):

from partial fractions? ln(y-1) - ln(y)

OpenStudy (fibonaccichick666):

we just agreed on the opposite I thought

OpenStudy (fibonaccichick666):

oh ignore that but still it's twoes

OpenStudy (fibonaccichick666):

not .5

OpenStudy (fibonaccichick666):

and oh god I need sleep

OpenStudy (fibonaccichick666):

yes you are correct

OpenStudy (fibonaccichick666):

100% correct

OpenStudy (anonymous):

haha sweet as, just gotta express it in terms of linear y now i guess. cheers, i reckon youve earnt some z's :)

OpenStudy (fibonaccichick666):

yes I get 0 then provided you did the calculation of ln|tanx+secx| right

OpenStudy (fibonaccichick666):

no, you just set it = to C

OpenStudy (fibonaccichick666):

for real, in differential eq. you don't solve for y anymore

OpenStudy (fibonaccichick666):

so the answer is 0= blah blah blah

OpenStudy (anonymous):

real talk, thanks for the help. espec considering the time of day haha

OpenStudy (fibonaccichick666):

but if you want, combine the logs nd then exp both sides

OpenStudy (anonymous):

yea s'what i did, just for aesthetics haha

OpenStudy (fibonaccichick666):

lnx-lny=ln(x/y) alrighty, night then! well morning I guess

OpenStudy (fibonaccichick666):

man I'm scatterbrained. Have fun with your dif!

OpenStudy (anonymous):

yea thanks enjoy your hard earned rest!